# A machine-learning photometric classifier for massive stars in nearby galaxies

### v1

@@ 2.3 Removing foreground stars
Table 2: Number of sources per photometric survey and foreground selection criteria per galaxy. The columns contain the initial number of Spitzer sources as well as their cross-matches with Pan-STARRS, Gaia , UHS, and VHS. Parallax and proper motion limits for foreground detection are provided, before the final number of sources after removing duplicate sources.
WLM 14234 3331 399 0 3291 --0.08+0.99 0.26$\pm$1.20 --0.10$\pm$1.11 14091 13139
NGC 55 8698 0 729 0 0 --0.06+1.13 0.02$\pm$1.20 0.06$\pm$1.49 8524 8496
IC 10 32901 4547 1931 6006 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 31673 29499
M31 815811 410634 26332 387613 0 --0.16+1.12 0.09$\pm$1.15 --0.13$\pm$1.27 809142 809052
NGC 247 13398 2470 621 0 0 0.01+0.85 0.21$\pm$0.98 --0.03$\pm$1.14 13095 13095
NGC 253 8734 1578 522 0 0 0.06+1.20 --0.00$\pm$1.07 0.02$\pm$1.31 8409 8381
NGC 300 20511 0 1400 0 11480 0.00+1.20 0.14$\pm$1.12 --0.12$\pm$1.23 20161 20153
IC 1613 28371 10364 1229 0 0 --0.18+0.96 0.11$\pm$1.03 --0.01$\pm$1.04 28245 26396
M33 73206 52455 11049 47594 0 --0.09+1.09 0.11$\pm$1.12 0.04$\pm$1.17 71847 71847
Phoenix Dwarf 10831 0 499 0 1212 0.01+0.87 0.35$\pm$0.95 --0.19$\pm$0.95 10703 10021
NGC 1313 6156 6156 481 6156 6156 0.27+1.36* 0.06$\pm$1.41* 0.45$\pm$1.14* 5970 5970
NGC 2366 495 156 64 0 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 462 462
NGC 2403 16644 3735 1517 0 0 --0.01+1.43 0.26$\pm$1.64 --0.05$\pm$1.47 15936 15910
M81 28479 3894 1072 0 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 27895 27875
Sextans B 4914 1166 141 0 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 4852 4413
NGC 3109 9474 2988 1069 0 0 --0.09+0.93 --0.01$\pm$1.06 --0.05$\pm$1.18 8939 8935
NGC 3077 2617 271 90 0 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 2548 2548
Sextans A 2888 880 219 0 355 0.03+0.84 --0.20$\pm$0.98 --0.12$\pm$0.94 2848 2693
NGC 4214 1159 368 95 89 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 1149 1149
NGC 4736 10043 1248 349 657 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 9861 9861
NGC 4826 4659 480 149 306 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 4577 4575
M 83 15020 2422 1396 0 3877 --0.04+0.76 0.31$\pm$1.20 0.07$\pm$1.05 14132 14132
NGC 5253 721 0 119 0 187 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 622 622
NGC 6822 25599 18659 7061 0 15205 0.01+1.13 --0.08$\pm$1.05 --0.24$\pm$1.16 22483 22471
Pegasus DIG 11316 2234 251 0 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 11147 10530
NGC 7793 5535 887 353 0 0 0.04+1.47 0.09$\pm$1.32 -0.04$\pm$1.06 5433 5425
777 Spitzer data for galaxies IC 10, IC 1613, Pegasus DIG, Phoenix Dwarf, Sextans A and B, and WLM are derived from Boyer et al. (2015); M33, M81, NGC 2403, NGC 247, NGC 300, NGC 6822, and NGC 7793 from Khan et al. (2015); NGC 2366, NGC 253, NGC 4214, NGC 5253, and NGC 55 from Williams & Bonanos (2016); M31, M81, NGC 3077, NGC 4736, NGC 1313, and NGC 4826 from Khan (2017). \$*\$\$*\$footnotetext: For those galaxies the parallax and proper motion criteria derived from M31 were used, as there were not enough data to determine these quantities by using their data only.
@@ 2.4 Quality cuts
The application of the classifier was straightforward. After building a source catalog for each of our galaxies we parsed it to the classifier, which identified and built the necessary features, i.e. calculated the color terms $r-i$, $i-z$, $z-y$, $y-[3.6]$, $[3.6]-[4.5]$. We used the pre-trained models from Paper I to obtain the predictions for each galaxy [^8] For the application of the classifier check the GitHub page of the project (https://github.com/gmaravel/pc4mas), where a python notebook and the models can be found..
@@
The classifier provides a classification among the following classes (see Paper I for details): BSG - a rather loose class of early-type and hot stars, YSG - evolved yellow type stars, RSG - red supergiants, BeBR - B[e] supergiants, LBV - Luminous Blue Variables, WR - Wolf-Rayet stars, GAL - galaxies including AGNs and background QSOs (i.e. all extragalactic outliers). The probability per class is provided for each of the algorithms used (i.e. Support Vector Machines, Random Forest, Multi-Layer Perceptron; see Paper I for more details), which are then combined (with equal weighting) to provide the final set of probabilities. For each source the final (predicted) class corresponds to the one with the highest probability of the ensemble approach (see Paper I for details).
@@
In Paper I we investigated the number of correct vs. incorrect sources with probability. We found that the mean values for the correct and incorrect classifications (in the combined sample of M31 and M33) were $0.86\pm 0.01$ and $0.60\pm 0.03$, respectively. A cut at 0.86 would significantly limit the sources to be considered since the classifier performs slightly worse for lower metallicities than M31 and M33. Therefore, we opted to use a cut at 0.66, equivalent to the mean value $+3\sigma$ of the incorrect classification distribution, which gave us the opportunity to consider a larger sample of sources while keeping the fraction of incorrectly classified sources relatively low (c.f. Fig. 6 and 8 of Paper I).
@@
Additionally, band availability (i.e. missing data) significantly affects the performance and robustness of the prediction model, specifically the probability. Our tests on missing data imputation (c.f. Fig. 9 in Paper I) indicate that the classifier can effectively manage up to two missing features (equivalent to three missing bands, since each feature is a color index) with a loss of accuracy of approximately 10%. This resilience extends to three missing features, albeit with an increased accuracy loss of less than 20%. More missing values lead to an approximate 40% loss, hence making the classifier unreliable.
@@
Given that missing values are a common occurrence for our sources, we identified two scenarios for selecting objects for further processing. The more balanced approach involved selecting sources with a final probability greater than 0.66 and band completeness exceeding 0.6 (i.e. missing two features, three bands in total). Alternatively, a more relaxed option considered sources with a probability greater than 0.50 and band completeness exceeding 0.4 (i.e. missing three features and 4 bands). However, the second scenario introduces a higher rate of false positives, increasing noise, due to more misclassifications, in the final results. Therefore, we decided to proceed exclusively with the first approach. To ensure flexibility, our published catalogs include data for all available sources, allowing users to apply their own selection criteria as needed.
2.5 Collecting known sources from literature
@@
To better understand the strengths and the limitations of our classifier we needed to compare its predictions with previously classified sources. For this reason, we undertook the very demanding task to search the literature for all known sources found in our galaxy sample. We limited our search to sources with secured spectral classification, i.e. that had been obtained with spectroscopy and not from photometric or other (e.g. variability) criteria. We have collected the complete spectral samples - to the best of our knowledge - for all the 26 target galaxies, accounting to 5273 sources (from 83 different works). This number includes all massive stars known, including candidates, as well as another $\sim 330$ point sources (such as carbon stars, background galaxies, Hii regions, planetary nebulae, clusters). The numbers and corresponding references per galaxy can be found in Table 7 of the appendix. We note here that no spectral classifications were available for 5 galaxies (i.e. NGC 2366, NGC 3077, NGC 4214, NGC 4826, NGC 5253). In all cases we carefully checked for and removed duplicates, keeping the most recent and precise classifications. The current catalog serves as the most complete source of reference for spectral classifications for massive stars and candidates.
@@ 3 Results
In this section, we provide the results of the application of the machine-learning classifier to our galaxy sample, in the form of a full catalog, a table providing the statistics per class (per galaxy) based on the most secure predictions, as well as some indicative color-magnitude diagrams.
@@ 3.1 Catalog description
In Table 3 we provide the (first few lines) of the final catalog for all sources from all galaxies, comprising 1,147,650 sources and spanning 78 columns. In the Table we provide the source ID, Spitzer coordinates, Gaia DR3 ID, proper motion and parallax, Spitzer , PS1, and near-IR photometry and errors, previous classification (if available), probabilities per class for each method including the ensemble one, as well as final class, final probability, and band completeness.
Table 3: Final source catalog with predicted classifications for all galaxies.
ID R.A. (J2000) Dec. (J2000) Gaia_DR3_ID Final_Class Final_Prob Band_Compl
(deg) (deg)
WLM-2 0.52017 -15.44622 WR 0.602 1.0
WLM-3 0.52012 -15.40681 WR 0.647 0.4
WLM-4 0.52012 -15.36033 RSG 0.859 1.0
WLM-5 0.52012 -15.46578 WR 0.618 0.2
WLM-7 0.52012 -15.44331 WR 0.648 0.2
WLM-8 0.52012 -15.41617 WR 0.574 0.2
WLM-9 0.52008 -15.50714 WR 0.612 0.2
WLM-10 0.52008 -15.46256 WR 0.536 0.2
WLM-11 0.52008 -15.56622 WR 0.438 0.6
WLM-12 0.52008 -15.42581 WR 0.615 0.2
WLM-13 0.52008 -15.43997 WR 0.613 0.2
WLM-14 0.52008 -15.54617 WR 0.558 0.2
WLM-16 0.52008 -15.56514 RSG 0.396 0.2
WLM-17 0.52004 -15.53664 YSG 0.476 0.6
WLM-18 0.52004 -15.55456 WR 0.473 0.2
WLM-19 $0$15.40131 WR 0.430 0.6
WLM-21 $0$15.46006 WR 0.646 0.2
WLM-22 $0$15.60042 WR 0.615 0.2
WLM-24 0.51996 -15.45511 WR 0.616 0.2
WLM-25 0.51996 -15.35928 RSG 0.487 0.2
 [^9] This table is available in its entirety in the VizieR/CDS catalog tool. A portion is shown here for guidance regarding its form and content.
@@ 3.2 Populations
Table 4: Number of predictions per galaxy, after applying the selection criteria on probability and band completeness (from Sect. 2.4).
@@
For further consideration and exploration of the results we only kept sources that satisfied the quality criteria as defined in Section 2.4. This allowed us to exclude the most uncertain predictions from the classifier, either due to the low probability or the number of missing values (below the band completeness threshold we set). These limits remove the noise and allow us to investigate the results and assess the performance of the classifier. Table 4 presents the total number of source classifications, along with the number of predictions for the classes of RSG, YSG, BSG, BeBR, WR, LBV, and GAL, independently.
Out of the 26 galaxies, 5 of them (namely NGC 55, NGC 300, Phoenix Dwarf, NGC 1313, NGC 5253) have zero sources (hence are not presented in Table 4). This is due to a lack of PS1 data which leads to more missing values than what we accepted (less than 0.6). Since only the IR data is available for these galaxies, the predictions would be highly uncertain and - therefore - were excluded from further analysis.
Overall, we find the highest number of predictions to be for RSGs and WR stars. In particular, M31, M33, NGC 6822, IC 1613, IC10, and M81 have more than 1000 classified sources. However, we already know (from Paper I) that WR predictions suffer from many false positives. We further discuss the performance of the classifier in Section 4.1.
@@ 3.3 Color-Magnitude Diagrams
Figure 2: Example of CMDs ($z$ vs. $r-z$ and [4.5] vs. $[3.6]-[4.5]$) for M31 (top) and NGC 2403 (bottom) with predicted classifications. We plot all sources that satisfy the quality cuts imposed in Section 2.4, and their photometric values come from the original data (i.e. we do not plot sources whose values have been imputed during the application of the classifier). We use cyan circles for Blue Supergiants (BSG), blue diamonds for B[e] Supergiants (BeBR), gray triangles for galaxies (GAL), green pentagons for Luminous Blue Variables (LBV), red crosses for Red Supergiants (RSG), purple empty triangles for Wolf-Rayet stars (WR), and yellow filled bottom-sided for Yellow Supergiants (YSG). The total number of sources per class is provided next to the class in the legend.
@@
We explored the locations of the predicted populations in Color-Magnitude Diagrams (CMDs). Fig. 2 shows optical and mid-IR CMDs for two example galaxies, M31 (as the most populated galaxy) and NGC 2403. We note that we do not plot sources whose values have been imputed during the application of the classifier, but only sources with their original photometry. We use different symbols to indicate the predicted class for each source, and we show the total number of predictions per class in the legend.
@@
We find that the position of sources matches their predictions. In the $z$ vs. $r-z$ CMD we see BSG located on the left (of approximately $r-z\sim 0$) while YSG and RSG extend to redder colors as expected. The few BeBR and LBV predicted are located (consistently) close to the BSG and WR. The latter ones display a much broader distribution which is not real in many cases, as this class suffers the most from false positives. This is more striking in the case of M31, where we note the ”line” of WR that extends from the bulk of sources (with $r-z\sim 0$ mag) to a point of $r-z\sim-5$ mag (and of $z\sim 27.5$ mag). Most probably the classifier confuses these faint but blue sources with WR, because M31 is a nearby galaxy and we can observe fainter stellar populations. In contrast, we do not see this in NGC 2403.
@@
Regarding the [4.5] vs. $[3.6]-[4.5]$ CMD we notice that the majority of the RSGs lie around $[3.6]-[4.5]\sim 0$ mag in both M31 and NGC 2403, consistent with what we would expect. For M31 though there are sources with bluer colors. Although there are indeed RSGs with bluer colors (e.g. Bonanos et al. 2009; de Wit et al. 2024) there is confusion with other populations (due to the proximity of M31). There is definitely confusion of WRs with other populations (WR is the class with the higher false positive results). In NGC 2403 WRs extend to slightly redder colors which is consistent with what we would expect. YSGs are located around 0.0 for both galaxies. However, the two very bright points in NGC 2403 are most probably foreground sources. In this CMD, GAL seems to become more evident from the bulk of the stellar populations and they are distinctly separated with $[3.6]-[4.5]>0.5$ mag.
@@
Overall, the position of the sources matches their predicted classification. There is of course a fraction of sources that is misclassified. This is due to a number of physical and technical reasons, such as the fact that strict boundaries in these parameter spaces between some classes do not exist (e.g. BSG and WR, YSG, and RSG), the fact that there could still be some contaminant sources either because these have not been part of the classifier’s training, or because there could be photometric errors (e.g. crowding, aperture contamination, or remaining foreground sources). We further explore these in Section 4.1.

### v2

@@ 2.3 Removing foreground stars
Table 4 summarizes the initial number of Spitzer sources; the crossmatches available with Pan-STARRS, Gaia , UHS, and VHS; parallax and proper motion limits for foreground detection (from Gaia ); the number of sources selected based on these criteria, and the final number of sources after removing duplicates. This is the number of sources per galaxy analyzed in this work.
@@ 2.4 Quality cuts
The application of the classifier was straightforward. After building a source catalog for each of our galaxies, we parsed it to the classifier, which identified and built the necessary features, i.e., calculated the color terms rir-i, izi-z, zyz-y, y[3.6]y-[3.6], [3.6][4.5][3.6]-[4.5], including missing data imputation if needed. This is possible for two reasons: first, because the compiled catalogs were drawn from independently curated datasets that already applied quality cuts; second, no preprocessing (e.g., signal-to-noise filtering) is necessary at this stage, since any low-confidence or noisy source can be removed at post-processing, as presented in this section. We used the pretrained models from Paper I to obtain the predictions for each galaxy [^8] For the application of the classifier check the GitHub page of the project (https://github.com/gmaravel/pc4mas), where a python notebook and the models can be found..
@@
The classifier provides a classification among the following classes (see Paper I for details): BSG -blue supergiant, a rather loose class of early-type and hot stars; YSG - evolved yellow type stars; RSG - red supergiants; BeBR - B[e] supergiants; LBV - luminous blue variables; WR - Wolf-Rayet stars; GAL - point-like extragalactic sources (Active Galactic Nuclei (AGNs), Quasi-Stellar Objects (QSOs), compact galaxies), which enables the classifier to identify and remove background contaminants projected along the line of sight. The probability per class is provided for each of the algorithms used (i.e., support vector machines, random forest, multi-layer perceptron; see Paper I for more details), which are then combined (with equal weighting) to provide the final set of probabilities. For each source, the final (predicted) class corresponds to the one with the highest probability of the ensemble approach (see Paper I for details).
@@
In Paper I, we investigated the number of correct versus incorrect sources with probability. We found that the mean values for the correct and incorrect classifications (in the combined sample of M31 and M33) were $0.86\pm 0.01$ and $0.60\pm 0.03$, respectively. A cut at 0.86 significantly limits the sources to be considered, since the classifier performs slightly worse for lower metallicities than M31 and M33. Therefore, we opted to use a cut at 0.66, equivalent to the mean value $3\sigma$ of the incorrect classification distribution, which gave us the opportunity to consider a larger sample of sources while keeping the fraction of incorrectly classified sources relatively low (cf., Figs. 6 and 8 of Paper I).
@@
Additionally, band availability (i.e., missing data) significantly affects the performance and robustness of the prediction model, specifically probability. Our tests on missing data imputation (cf., Fig. 9 in Paper I) indicated that the classifier can effectively manage up to two missing features (equivalent to three missing bands, since each feature is a color index) with a loss of accuracy of approximately 10%. This resilience extends to three missing features, albeit with an increased accuracy loss of less than 20%. More missing values lead to an approximate 40% loss, making the classifier unreliable. The reported accuracy losses correspond to the overall classifier performance, although individual classes are affected differently (cf., Fig. 12 in Paper I). However, class-specific adjustments are not feasible during the application of the classifier since the true class is unknown.
@@
Given that missing values are a common occurrence for our sources, we identified two scenarios to select objects for further processing. The more balanced approach involved selecting sources with a final probability greater than 0.66 and band completeness exceeding 0.6 (i.e., missing two features, three bands in total). Alternatively, a more relaxed option considered sources with a probability greater than 0.50 and band completeness exceeding 0.4 (i.e., missing three features and four bands). However, the second scenario introduced a higher rate of false positives, increasing noise, due to more misclassifications, in the final results. Therefore, we decided to proceed exclusively with the first approach. To ensure flexibility, our published catalogs include data for all available sources, allowing users to apply their own selection criteria as needed.
2.5 Collecting known sources from the literature
@@
To better understand the strengths and limitations of our classifier, we needed to compare its predictions with previously classified sources. For this reason, we undertook the very demanding task of searching the literature for all known sources found in our galaxy sample. This task complements our work presented in Paper I, where, during the development of the classifier, we were faced with many constraints imposed by Gaia DR2 coverage and smaller galaxy samples. We limited our search to sources with secured spectral classification, i.e., that had been obtained with spectroscopy and not from photometric or other (e.g., variability) criteria. We collected complete spectral samples - to the best of our knowledge - for all the 26 target galaxies, accounting for 5273 sources (from 83 different works). This number includes all known massive stars, including candidates, as well as another $\sim 330$ point sources (such as carbon stars, background galaxies, Hii regions, planetary nebulae, and clusters). The numbers and corresponding references for each galaxy can be found in Table 7. We note here that no spectral classifications were available for five galaxies (i.e., NGC 2366, NGC 3077, NGC 4214, NGC 4826, and NGC 5253). In all cases, we carefully checked for and removed duplicates, keeping the most recent and precise classifications. The current catalog serves as the most complete source of reference for spectral classifications for massive stars and candidates beyond the Milky Way and the Clouds.
@@ 3 Results
In this section, we provide the results of the application of the machine-learning classifier to our galaxy sample, in the form of a full catalog. This table contains the statistics per class (per galaxy) based on the most secure predictions. We also present and discuss some indicative CMDs derived directly from these results.
@@ 3.1 Catalog description
In Table 5 we provide the first few lines of the final catalog for all sources from all galaxies, comprising 1,147,650 sources and spanning 78 columns. In the Table we provide the source ID number, Spitzer coordinates, Gaia DR3 ID number, proper motion and parallax, Spitzer , PS1, and near-IR photometry and errors, previous classification (if available), probabilities per class for each method including the ensemble one, as well as final class, final probability, and band completeness.
@@ 3.2 Populations
Table 2: Number of predictions per galaxy, after applying selection criteria on probability and band completeness (from Sect. 2.4).
@@
For further consideration and exploration of the results, we only kept sources that satisfied the quality criteria as defined in Sect. 2.4. This excluded the most uncertain predictions from the classifier which add noise, either due to the low probability or the number of missing values (below the band completeness threshold we set). Consequently, it allowed us to investigate the results, and assess the performance of the classifier. Table 2 presents the total number of source classifications, along with the number of predictions for the classes of RSG, YSG, BSG, BeBR, WR, LBV, and GAL, independently.
Of the 26 galaxies, five (namely NGC 55, NGC 300, Phoenix Dwarf, NGC 1313, and NGC 5253) have zero sources (hence are not presented in Table 2). This is due to a lack of PS1 data, which leads to more missing values than we accepted (less than 0.6). Since only the IR data are available for these galaxies, the predictions were highly uncertain and were therefore excluded from further analysis.
Overall, we find that the highest number of predictions is for RSGs and WR stars. In particular, M31, M33, NGC 6822, IC 1613, IC10, and M81 have more than 1000 classified sources. However, we already know (from Paper I) that WR predictions suffer from many false positives. We further discuss the performance of the classifier in Sect. 4.1.
@@ 3.3 Color-magnitude diagrams
Figure 2: Example of CMDs ($z$ vs. rzr-z and [4.5] vs. [3.6][4.5][3.6]-[4.5]) for M31 (top) and NGC 2403 (bottom) with predicted classifications. We plot all sources that satisfy the quality cuts imposed in Sect. 2.4, and their photometric values come from the original data (i.e., we do not plot sources whose values have been imputed during the application of the classifier). We use cyan circles for blue supergiants (BSG), blue diamonds for B[e] supergiants (BeBR), gray triangles for galaxies (GAL), green pentagons for luminous blue variables (LBV), red crosses for red supergiants (RSG), purple empty triangles for Wolf-Rayet stars (WR), and yellow filled bottom-sided for yellow supergiants (YSG). The total number of sources per class is provided next to the class in the legend.
@@
We explored the locations of the predicted populations in CMDs. Figure 2 shows optical and mid-IR CMDs for two example galaxies, M31 (as the most populated galaxy) and NGC 2403. We note that we do not plot sources whose values have been imputed during the application of the classifier, but only sources with their original photometry. We use different symbols to indicate the predicted class for each source and show the total number of predictions per class in the legend.
@@
We find that the positions of the sources match their predictions. In the $z$ versus rzr-z CMD we see BSG located on the left (of approximately $r-z\sim 0$) while YSG and RSG extend to redder colors as expected. The few predicted BeBR and LBV are (consistently) located close to the BSG and WR. The latter display a much broader distribution, which is not real in many cases, as this class suffers the most from false positives. This is more striking in the case of M31, where we note a ”tail” of WR that extends from the bulk of sources (with $r-z\sim 0$ mag) to a point of $r-z\sim-5$ mag (and of $z\sim 27.5$ mag). We attribute this trend to a combination of faint-end detection bias and color-space overlap between the WR and other classes (including additional faint blue populations visible at the distance of M31, for which the classifier is not trained), rather than from systematically noisy photometry. In contrast, we do not see this in NGC 2403.
@@
Regarding the [4.5] versus [3.6][4.5][3.6]-[4.5] CMD we notice that the majority of RSGs lie around $3.6]-[4.5]\sim 0$ mag in both M31 and NGC 2403, consistent with what we would expect. For M31 though, there are sources with bluer colors. Although there are indeed RSGs with bluer colors (e.g., Bonanos et al. 2009; de Wit et al. 2024), there is confusion with other populations (due to the proximity of M31). There is definitely confusion of WRs with other populations (WR is the class with the highest false positive results). In NGC 2403 WRs extend to slightly redder colors, which is consistent with what we expect. Yellow supergiants are located around 0.0 for both galaxies. However, the two very bright points in NGC 2403 (sources NGC2433-37 and NGC2433-66) are probably foreground sources, as there are no Gaia data that can confirm their nature. In this CMD, GAL seems to become more evident from the bulk of the stellar populations, and they are distinctly separated with [3.6][4.5]>0.5[3.6]-[4.5]>0.5 mag.
@@
Overall, the position of the sources matches their predicted classification. Of course, there are a fraction of sources that are misclassified. This is due to a number of physical and technical reasons, such as the fact that strict boundaries in these parameter spaces between some classes do not exist (e.g., BSG and WR, YSG, and RSG), the fact that there could still be some contaminant sources, either because these have not been part of the classifier’s training or because there could be photometric errors (e.g., crowding, aperture contamination, or remaining foreground sources). We further explore these in Sect. 4.1.